OpenStudy (anonymous):

MEEDAAL! What is the measure of the exterior angle x? 58° 80° 100° 180°

OpenStudy (anonymous):



OpenStudy (anonymous):

@ajprincess @amistre64 @AngelicDragon @experimentX @radar @LazyBoy

OpenStudy (anonymous):

@satellite73 @UnkleRhaukus

OpenStudy (anonymous):

@ganeshie8

OpenStudy (aravindg):

Sum of the interior angles should be 360 Y+90+112+58=360 Solve for Y Once you get Y see that X and Y form a linear pair Hence X+Y=180 Solve for X

OpenStudy (anonymous):

100?

OpenStudy (anonymous):

x = 100

OpenStudy (jhannybean):

yep. 360-260=100.

OpenStudy (magbak):

x= 100